OpenStudy (anonymous):

how to solve it? 4x+y=6 and y=6-4x?am i going to use the substitution or addition method?

OpenStudy (anonymous):

this is impossible if the numbers X and Y are not natural so the only solution is x=1 and y=2 if they are both natural numbers but if not then we have a lot of solutions.

OpenStudy (anonymous):

using the linear equation method like addition and substitution method? are you sure?

OpenStudy (anonymous):

y=6-4x 6-4x=6-4x 0=0

OpenStudy (anonymous):

but here there is only one equation and from one equation with 2 unknowns the only solution is to try substituting random numbers until you find one solution but that happens only if they are natural numbers otherwise you can never guess what the solutions are because there will be many

OpenStudy (anonymous):

tnx..=)

OpenStudy (anonymous):

np :)